OpenStudy (raphaelll):

Rearrange the following formula to make y the subject of the equation. a= 3y-4 ______ 6

jimthompson5910 (jim_thompson5910):

I'll show you how to solve a similar problem (just with different numbers) \[\Large a = \frac{8y - 5}{2}\] \[\Large {\color{red}{2*}}a = {\color{red}{2*}}\frac{8y - 5}{2} \ \text{multiply both sides by 2}\] \[\Large 2a = 8y - 5\] \[\Large 2a{\color{red}{+5}} = 8y - 5{\color{red}{+5}} \ \text{add 5 to both sides}\] \[\Large 2a+5 = 8y\] \[\Large \frac{2a+5}{{\color{red}{8}}} = \frac{8y}{{\color{red}{8}}} \ \text{divide both sides by 8}\] \[\Large \frac{2a+5}{8} = y\] \[\Large y = \frac{2a+5}{8}\] again, this is not the answer as it's just a similar example

OpenStudy (raphaelll):

so would the answer be y = 2a+4 _______ 3